OpenStudy (anonymous):

Hybrid vehicle: Purchase price=$29,500 Mileage rating= 28miles per gallon Tax credit=$2,200 Non hybrid: Purchase price=$25,000 Mileage rating=20miles per gallon Tax credit=none. If the average price of the gas will be $1.95 per gallon, how many miles of driving will it take before fuel savings of the hybrid will make up for it greater initial lost.

OpenStudy (jiteshmeghwal9):

Try to mention any one to call like @ash2326 @amistre64 @satellite73 @jhonyy9 :)

OpenStudy (jiteshmeghwal9):

mention any one for help:)

OpenStudy (anonymous):

how to do that? just send message to them?

OpenStudy (ash2326):

nope, just @ followed by the user name. like this @nia_maepenix BTW I'll help you with this

OpenStudy (anonymous):

thanks a lot. I'm kinda new here

OpenStudy (ash2326):

Let x be the no. of miles traveled, at which the hybrid fuel savings make up for the greater cost

OpenStudy (anonymous):

yes

OpenStudy (theviper):

from whom do u want help type there name & just add @ before there name ;)

OpenStudy (anonymous):

the cost of the hybrid is \(\$29,500- \$2,200=\$27,300\) and the cost of the non-hybrid is \(\$25,000\) so the hybrid cost \(\$27,300-\$25,000=\$2,300\) more

OpenStudy (anonymous):

cost-tax=?

OpenStudy (anonymous):

tax CREDIT

OpenStudy (anonymous):

so the answer is2,300?the lost

OpenStudy (anonymous):

@ash2326 can u finish your explanation too?

OpenStudy (anonymous):

the \($2,300\) is how much more the hybrid cost we are not done yet by any means

OpenStudy (anonymous):

thanks satelite73

OpenStudy (anonymous):

we have to figure out how many miles we have to go before the fuel savings makes up for the extra \($2,300\) we spent on the hybrid

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so 2300/1.95?

OpenStudy (anonymous):

oh no hold on a sec

OpenStudy (anonymous):

ok... thanks a lot

OpenStudy (anonymous):

lets imagine we go 280 miles in the hybrid we would use \(280\div28=10\) gallons of gas at a cost of \(10\times 1.95=19.5\) dollars in the non hybrid we would use \(280\div 20=14\) gallons of gas at a cost of \(14\times 1.95=27.3\) dollars which is obviously more expensive. by how much? by \(27.3-19.5=7.80\)

OpenStudy (anonymous):

so we need to have exact miles so we can figure out how much is lost?

OpenStudy (anonymous):

we want to know when this difference will equal 2,300 so we need to introduce a variable lets say \(x\) is the number of miles we drive then in the hybrid will we use \(\frac{x}{28}\) gallons of gas at a cost of \(1.95\times \frac{x}{28}\) and in the non hybrid we will use \(\frac{x}{20}\) gallons of gas at a cost of \(1.95\frac{x}{20}\)

OpenStudy (anonymous):

hybrid is 1.95/28x and non is 1.95/20x

OpenStudy (anonymous):

our job is to solve \[1.95\times \frac{x}{20}-19.5\times \frac{x}{28}=2,300\]

OpenStudy (anonymous):

typo there, should be \[1.95\times \frac{x}{20}-1.95\times \frac{x}{28}=2,300\]

OpenStudy (anonymous):

we can do it as follows \[1.95(\frac{x}{20}-\frac{x}{28})=2,300\] \[1.95(\frac{28x-20x}{20\times 28})=2,300\] \[\frac{8x}{560}=2,300\div 1.95=1179.5\] rounded \[\frac{x}{70}=1179.5\] \[x=1179.5\times 70\]

OpenStudy (anonymous):

so the lost is 82565?

OpenStudy (anonymous):

i don't know i didn't multiply

OpenStudy (anonymous):

i think divide

OpenStudy (anonymous):

1179.5/70

OpenStudy (anonymous):

but before i do i hope it is clear what i did i was not sure what equation i needed to solve, so i did it with numbers to see how it worked that is why i wrote the first post about driving 280 miles

OpenStudy (anonymous):

multiply don't divide

OpenStudy (anonymous):

yeah i get it

OpenStudy (anonymous):

you have \(\frac{x}{70}=1179.5\) so \(x=1179.5\times 70=82565\)

OpenStudy (anonymous):

let me check to see if this looks right

OpenStudy (anonymous):

thanks. this is the lost?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

miles of driving will it take before fuel savings of the hybrid will make up for it greater initial lost is 82565